OpenStudy (anonymous):

A trapezoid is drawn on a grid with vertices at P (-5, 7), Q (-2, 7), R (2, 3), and S (-8, 3). Trapezoid UVWX will be drawn similar to Trapezoid PQRS with side WX corresponding to side RS. The point (1, 6) will be the location of vertex W and point (-4, 6) will be the location of point X. What will be the coordinates of vertex V? Show all your work or explain your reasoning.

OpenStudy (albert0898):

WX corresponds to the long side of P Q R S RS = 10 , WX = 5 so the dimensions of the new trapezoid are half the original. x-distance P-S = 3 so x-distance U-X = 3/2 and the x-coordinate of U is -4+3/2 = -5/2 x-distance Q-R = 4 so x-distance W-V = 2 and the x-coordinate of V is 1-2 = -1 They probably intend for U-V to be above W-X, in which case the y-coordinates of U and V would be 6+2 = 8 But it could just as well be below, in which case the y-coordinates would 6 - 2 = 4 In summary, U-V above X-W U = (-5/2, 8) V = (-1,8) [ probably the intended answer ] or U-V below X-W U = (-5/2, 4) V = (-1, 4) [ an equally valid answer ] Similar figures do not have to be in the same orientation. NOTE: Corresponding means similar in function or in character!

OpenStudy (anonymous):

which one should I put though? the above or below?